# Information Discovery in e-Commerce

### v1

@@ 6.3.4 Retraining strategies
E-commerce recommender systems rely on knowledge gleaned from historical interactions. As the collected interactions grows, recommendation models must be regularly retrained to reflect users’ dynamic preferences. There are two intuitive heuristic retraining methods:
@@
Numerous deep learning-based retraining methods have been proposed. SML proposed a sequential meta learning model (SML), including a meta-learning retraining framework for vanilla matrix factorization models. SML model captures the trend of changes between two distinct retraining phases at adjacent times. Recently, graph convolutional neural network (GCN) has become the cutting-edge technique for recommendation [LightGCN, pinsage]. However, GCN-based recommender models encounter challenges regarding model retraining as GCN-based models take more time to converge. CIGC [CIGC] has been proposed with two novel operators: incremental graph convolution and colliding effect distillation. The incremental graph convolution estimates the output of fully retraining the graph convolution using only new data; whereas the colliding effect distillation uses causal inference to retrain the representations of users (or items) that have no new data.
@@ 6.4 Re-ranking strategies
The two-stage framework still faces several problems in e-commerce recommendation: (i) The point-wise objective functions (e.g., log-loss) at the ranking stage often become sub-optimal because of the neglect of mutual influences between items. (ii) Users with different preferences may exhibit different behavior patterns. (iii) Recommendation result diversification has not yet been well addressed during the ranking stage. Therefore, before exposing the recommendation items to users, most of e-commerce recommender systems usually refine the recommendation results through an additional re-ranking stages. Generally, the goal of re-ranking is to enhance the recommendation results through additional criteria or constraints [Chen2017Improving, zehlike2017fa, abdollahpouri2019popularity]. Existing methods on re-ranking can be categorized into two types: heuristic strategies and list-wise objective functions. The former type of methods are based on determinantal point processes [Wilhelm2018Practical, Chen2017Improving] and maximal marginal relevance [Carbonell1998The]. These methods have been widely applied in e-commerce to avoid the items with the same category being presented consecutively for greater diversity. While for list-wise objective functions, DLCM [ai2018learning] and PRM [Pei2019Personalized] were proposed with specific list-wise optimization objectives.
@@ 6.5 Emerging directions
We discuss five emerging research directions on e-commerce recommendation: (i) structured recommendations, (ii) conversational recommendation, (iii) reasoning recommendations and explanations, (iv) biases and debiasing in recommendation, and (v) unifying recommendation and search.
@@ 6.5.1 Structured recommendations
In applications like music or bundle recommendations, the objective is to provide users with a ’slate’—a combination of items—to maximize their engagement with the recommended content. This task raises critical questions, including the consideration of metrics such as diversity and the computational challenges posed by the combinatorial nature of slates. Reinforcement learning (RL) is extensively applied in slate recommendation[Ie2019ReinforcementLF, Sunehag2015DeepRL, DBLP:conf/kdd/TomasiCKCRD23]. However, due to the combinatorial complexity of actions, RL typically necessitates simplifying assumptions, such as the user selecting the optimal item [Ie2019ReinforcementLF]. An alternative approach involves integrating a separate user preference model to optimize slate assembly and subsequently training the RL model [DBLP:conf/kdd/TomasiCKCRD23]. DBLP:conf/nips/SwaminathanKADL17 investigated off-policy evaluation and optimization via inverse propensity scores for slate interactions. DBLP:conf/www/MehrotraLKLH19 constructed a hierarchical model to assess user satisfaction in slate recommendation systems.
@@
The music playlist recommendation can be considered as a special case of music recommendation, focusing on delivering a curated list of songs to users. The order and characteristics of music tracks significantly influence the playlist’s overall quality. An earlier study employed time-series-based machine learning to address the challenge of recommending music playlists[Choi2016TowardsPG, Irene2019AutomaticPG, Monti2018AnEA, Vall2018TheIO, Kim2018TowardsSM, DBLP:conf/dasfaa/YangZXYZZ19]. Choi2016TowardsPG utilized a recurrent neural network (RNN) for music playlist generation, emphasizing track transition qualities. Monti2018AnEA implemented an ensemble of RNNs, leveraging pre-trained embeddings for album and title representation. Irene2019AutomaticPG predicted user preferences by analyzing manually created playlists, employing both RNN and convolutional neural network (CNN) models. Recent studies have employed reinforcement learning (RL)-based methods to capture users’ long-term preferences[Hu2017PlaylistRB, DBLP:conf/ismir/ShihC18, DBLP:conf/gcce/SakuraiTOH20, DBLP:conf/lifetech/SakuraiTOH21, DBLP:journals/sensors/SakuraiTOH22, DBLP:conf/kdd/TomasiCKCRD23, DBLP:conf/atal/LiebmanSS15]. DBLP:conf/atal/LiebmanSS15 attempted to utilize a novel reinforcement-learning-based music recommendation system that generates playlists by considering both song preferences and transitions. Hu2017PlaylistRB enhanced playlist generation performance by integrating user feedback into the recommendation reward function. DBLP:conf/ismir/ShihC18 incorporated novelty and popularity indices into the reward function, resulting in playlists with a mix of new and well-known tracks. DBLP:journals/sensors/SakuraiTOH22 utilized informative knowledge graphs to enhance reinforcement learning optimization, and allowing users to customize flexible reward functions to discover new music genres. DBLP:conf/kdd/TomasiCKCRD23 presented a reinforcement learning framework optimizing directly for user satisfaction via the use of a simulated playlist-generation environment.
@@
Next-basket recommendation(NBR) is a task focused on predicting a user’s next shopping basket based on their past purchase history, aiming to enhance user experience and satisfaction. There are mainly three families of NBR methods. First are conventional NBR methods, such as those employing pattern mining[DBLP:conf/icdm/GuidottiRPGP17], KNN models[DBLP:conf/sigir/Hu0GZ20, DBLP:conf/um/FaggioliPA20], and Markov chain models[DBLP:conf/www/RendleFS10]. DBLP:conf/sigir/Hu0GZ20 and DBLP:conf/um/FaggioliPA20 model temporal patterns across frequency data and integrate this with neighbor information or user-wise collaborative filtering. DBLP:conf/www/RendleFS10 leverage matrix factorization and Markov chains to model users’ general interest and basket transition relations. Second are latent representation methods, which utilized representation learning techniques to capture implicit patterns in data. For instance, DBLP:conf/sigir/WangGLXWC15 applied aggregation operations to learn a hierarchical representation of user’s last basket to predict the next basket. Third are deep learning-based method. Recurrent neural networks (RNNs) have been extensively applied in next-basket recommendation, demonstrating their efficacy in learning long-term trends by modeling the whole basket sequence. For instance, DBLP:conf/sigir/YuLWWT16 utilized max/avg pooling to encode baskets and DBLP:conf/kdd/HuH19 adapted an attention mechanism and integrate frequency information to improve performance. Some methodsDBLP:conf/ijcai/LeLF19, DBLP:conf/aaai/WangH0SOC20 utilized item relations to enhance representation. DBLP:conf/kdd/YuSDL0L20 employed graph neural networks (GNNs) to model item-item relations between baskets and a self-attention mechanism to discern temporal dynamics. Some methods [DBLP:conf/sigir/BaiNZZDW18, DBLP:journals/corr/abs-2109-11654, DBLP:conf/dasfaa/LengYXX20, DBLP:conf/sigir/SunBDL0L20, DBLP:journals/jcst/WangZNG19] leverage auxiliary information, including product categories, amounts, prices, and explicit timestamps.
@@ 6.5.2 Conversational recommendation
The task of a conversational recommender system (CRS) is to provide recommendations to users through conversational interactions. Since recently, CRSs are attracting increasing attention [zhao2013interactive, ConversationRS-KDD16, yu2019visual, zou2020neural, mangili2020bayesian, ConversationRS-SIGIR18, lei20estimation, lei2020interactive, zhou2020improving, liu2020towards, zhou2020topicguided, zhang2020conversational, li2020seamlessly]. According to [gao-2021-advances], the task of CRS is formally defined as follows:
@@
Building on advances in interactive recommendation [christakopoulou2018q, wang2017factorization, liu2020diversified], early studies on CRSs formulated the task as a specific application of task-oriented multi-turn dialogue systems (TDS) [DBLP:conf/ccks/LeLWWLJ18, dhingra-EtAl:2017:Long1, wen2017network, zhang2019memory]. For existing studies on CRSs, there are two main types of strategies: attribute-aware and topic-guided.
@@
Attribute-aware CRSs aim to answer three main research questions: “whether to ask or recommend”, “which attributes to ask” or “which items to recommend.” Early work on attribute-aware CRSs obtained user preferences based on asking about items directly [zhao2013interactive, wang2018online, ConversationRS-KDD16, zou2020neural, DBLP:conf/aaai/VendrovLHB20], or asking attributes through a heuristic method [christakopoulou2018q, zhang2018towards, DBLP:conf/sigir/LuoYWS20]. There are two main kinds of attribute-aware CRSs solutions. One kind asks a fixed number of questions and makes a recommendation at the last turn [lei20estimation, lei2020interactive]; whereas the other predicts a specific turn to recommend items. Reinforcement learning strategies have been successfully applied to attribute-aware CRSs. liu2020towards and li2020seamlessly focused on cold-start users in conversational recommendation and extended bandit-based algorithms to balance the trade-off between exploration and exploitation. zou2022improving proposed TSCR, a Transformer-based sequential conversational recommendation method that captures the sequential dependencies in dialogues to enhance recommendation accuracy. deng2023unified propose a novel Unified MultI-goal conversational recommender system, namely UniMIND, which unifies these goals into a single sequence-to-sequence (Seq2Seq) paradigm and employs prompt-based learning strategies to facilitate multi-task learning.
@@
Topic-guided CRSs interact with users through natural language conversations with fluent responses and precise recommendations [DBLP:conf/nips/LiKSMCP18, zhou2020topicguided, DBLP:conf/emnlp/ChenLZDCYT19, liu2020towards, zhou2020improving, DBLP:conf/emnlp/MaTH21, DBLP:conf/wsdm/ZhouZZWJ022]. Unlike attribute-aware CRSs, topic-guided CRSs focus on making recommendations using free text, which creates considerable flexibility to influence how a dialogue continues. External knowledge has been applied in topic-guided CRSs [DBLP:conf/emnlp/MaTH21, zhou2020improving, DBLP:conf/emnlp/ChenLZDCYT19]. DBLP:conf/emnlp/ChenLZDCYT19 integrated a recommendation system and a dialogue system via an end-to-end framework to bridge the gap between the two systems. DBLP:conf/nips/LiKSMCP18 used an auto-encoder for recommendation and a hierarchical RNN for response generation. zhou2020topicguided proposed a topic-guided CRSs method that incorporates topic threads to enforce transitions actively toward a final recommendation. Recently, external knowledge graphs have been shown to be effective in improving the performance of topic-guided conversational recommendation systems. DBLP:conf/emnlp/ChenLZDCYT19 applied knowledge graphs to enhance the semantics of contextual items for recommendation. zhou2020improving incorporated both word-oriented and entity-oriented knowledge graphs. DBLP:conf/emnlp/MaTH21 performed a tree-structured reasoning on a knowledge graph for recommendation. zhang2022analyzing focused on user reformulation behaviors to improve the robustness of conversational agents. ren2022variational explored user preferences in conversational recommendation, proposed a variational reasoning mechanism to jointly track both short-term and long-term user behaviors. zhang2023variational made the first attempt to explicitly address the problem of dynamic reasoning over incomplete knowledge graphs. However, none of the existing studies are capable of fusing recommendation and response generation in an end-to-end manner, which limits the potential for mutual reinforcement between these two tasks. Additionally, a lack of interpretability in current conversational recommendation system (CRS) models further hinders their ability to fully align with user needs. Existing models are typically trained on conversational recommendation datasets, but the assumption that the standard items and responses in these benchmark datasets are optimal leads to a tendency for CRSs to replicate the logic of the recommenders found in the data, rather than truly addressing the evolving needs of the users. This misalignment remains a significant challenge in advancing more user-centric and adaptable conversational recommendation systems.
@@
Although CRSs poses many merits, its evaluation is still a thorny issue. Recent studies have evaluated CRSs either through offline evaluation or human evaluation [Lamel2000TheLA, li2015toward]. Offline evaluation evaluates a dialogue system based on test sets, whereas human evaluation reflects the overall performance of the agent through in-field experiments [Black2011SpokenDC, gilotte2018offline] or crowd-sourcing [zhou2020topicguided, DBLP:conf/nips/LiKSMCP18]. However, offline evaluation is often limited to single turn assessments, while human evaluation is intrusive, time-intensive, and is not scalable [zhao2019toward]. As an alternative, user simulators that mimic user behavior are able to provide broad insights to generate human-like conversations for assessing CRSs [afzali2023usersimcrs].

### v3

@@ 6.3.4 Retraining strategies
E-commerce recommender systems rely on knowledge gleaned from historical interactions. As the number of collected interactions grows, recommendation models must be regularly retrained to reflect users’ dynamic preferences. There are two intuitive heuristic retraining methods:
@@
Numerous deep learning-based retraining methods have been proposed. SML propose a sequential meta-learning model (SML), including a meta-learning retraining framework for vanilla matrix factorization models. SML model captures the trend of changes between two distinct retraining phases at adjacent times. Recently, graph convolutional neural network (GCN) has become the cutting-edge technique for recommendation [LightGCN, ying2018graph]. However, GCN-based recommender models encounter challenges regarding model retraining as GCN-based models take more time to converge. CIGC [CIGC] has been proposed with two novel operators: incremental graph convolution and colliding effect distillation. The incremental graph convolution estimates the output of fully retraining the graph convolution using only new data; whereas the colliding effect distillation uses causal inference to retrain the representations of users (or items) that have no new data.
@@ 6.4 Re-ranking strategies
The two-stage recommendation framework faces several problems in e-commerce recommendation: (i) The point-wise objective functions (e.g., log-loss) at the ranking stage often become sub-optimal because of the neglect of mutual influences between items. (ii) Users with different preferences may exhibit different behavior patterns. (iii) Recommendation result diversification has not yet been well addressed during the ranking stage. Therefore, before exposing recommended items to users, most e-commerce recommender systems refine the results through an additional re-ranking stage. Generally, the goal of re-ranking is to enhance the recommendation results through additional criteria or constraints [Chen2017Improving, zehlike2017fa, abdollahpouri2019popularity]. Re-ranking methods can be categorized into heuristic strategies and list-wise objective functions. The former type of method is based on determinantal point processes [Wilhelm2018Practical, Chen2017Improving] and maximal marginal relevance [Carbonell1998The]. These methods have been widely applied in e-commerce to avoid the items with the same category being presented consecutively for greater diversity. While for list-wise objective functions, DLCM [ai2018learning] and PRM [Pei2019Personalized] have been proposed with specific list-wise optimization objectives.
@@ 6.5 Emerging directions
We discuss five emerging research directions in e-commerce recommendation: (i) structured recommendations, (ii) conversational recommendation, (iii) reasoning recommendations and explanations, (iv) biases and debiasing in recommendation, and (v) unifying recommendation and search.
@@ 6.5.1 Structured recommendations
In applications like music or bundle recommendations, the objective is to provide users with a “slate” – a combination of items – to maximize their engagement with the recommended content. This task raises critical questions, including the consideration of metrics such as diversity and the computational challenges posed by the combinatorial nature of slates. Reinforcement learning (RL) is extensively applied in slate recommendation [Ie2019ReinforcementLF, Sunehag2015DeepRL, deffayet-2023-generative, DBLP:conf/kdd/TomasiCKCRD23]. However, due to the combinatorial complexity of actions, RL typically necessitates simplifying assumptions, such as the user selecting the optimal item [Ie2019ReinforcementLF]. An alternative approach involves integrating a separate user preference model to optimize slate assembly and subsequently training the RL model [DBLP:conf/kdd/TomasiCKCRD23]. DBLP:conf/nips/SwaminathanKADL17 investigate off-policy evaluation and optimization via inverse propensity scores for slate interactions. DBLP:conf/www/MehrotraLKLH19 constructe a hierarchical model to assess user satisfaction in slate recommendation systems.
@@
Music playlist recommendation can be considered as a special case of music recommendation, focusing on delivering a curated list of songs to users. The order and characteristics of music tracks significantly influence the playlist’s overall quality. An earlier study employs time-series-based machine learning to address the challenge of recommending music playlists [Choi2016TowardsPG, Irene2019AutomaticPG, Monti2018AnEA, Vall2018TheIO, Kim2018TowardsSM, DBLP:conf/dasfaa/YangZXYZZ19]. Choi2016TowardsPG use a recurrent neural network (RNN) for music playlist generation, emphasizing track transition qualities. Monti2018AnEA implement an ensemble of RNNs, using pre-trained embeddings for album and title representation. Irene2019AutomaticPG predict user preferences by analyzing manually created playlists, employing both RNN and convolutional neural network (CNN) models. Later studies have employed reinforcement learning (RL)-based methods to capture users’ long-term preferences [Hu2017PlaylistRB, DBLP:conf/ismir/ShihC18, DBLP:conf/gcce/SakuraiTOH20, DBLP:conf/lifetech/SakuraiTOH21, DBLP:journals/sensors/SakuraiTOH22, DBLP:conf/kdd/TomasiCKCRD23, DBLP:conf/atal/LiebmanSS15]. DBLP:conf/atal/LiebmanSS15 use a novel reinforcement-learning-based music recommendation system that generates playlists by considering both song preferences and transitions. Hu2017PlaylistRB enhance playlist generation performance by integrating user feedback into the recommendation reward function. DBLP:conf/ismir/ShihC18 incorporate novelty and popularity indices into the reward function, resulting in playlists with a mix of new and well-known tracks. DBLP:journals/sensors/SakuraiTOH22 use informative knowledge graphs to enhance reinforcement learning optimization, and allowing users to customize flexible reward functions to discover new music genres. DBLP:conf/kdd/TomasiCKCRD23 present a reinforcement learning framework optimizing directly for user satisfaction via the use of a simulated playlist-generation environment.
@@
Next-basket recommendation (NBR) is a task focused on predicting a user’s next shopping basket based on their past purchase history, aiming to enhance user experience and satisfaction. There are mainly three families of NBR methods. First are conventional NBR methods, such as those employing pattern mining [DBLP:conf/icdm/GuidottiRPGP17], KNN models [hu-2020-modeling, DBLP:conf/um/FaggioliPA20], and Markov chain models [FPMC]. hu-2020-modeling and DBLP:conf/um/FaggioliPA20 model temporal patterns across frequency data and integrate this with neighbor information or user-wise collaborative filtering. FPMC use matrix factorization and Markov chains to model users’ general interest and basket transition relations. Second are latent representation methods, which use representation learning techniques to capture implicit patterns in data. For instance, DBLP:conf/sigir/WangGLXWC15 apply aggregation operations to learn a hierarchical representation of user’s last basket to predict the next basket. Third are deep learning-based method. Recurrent neural networks (RNNs) have been extensively applied in next-basket recommendation, demonstrating their efficacy in learning long-term trends by modeling the whole basket sequence. For instance, DBLP:conf/sigir/YuLWWT16 use max/avg pooling to encode baskets and DBLP:conf/kdd/HuH19 adapt an attention mechanism and integrate frequency information to improve performance. Some methods [DBLP:conf/ijcai/LeLF19, DBLP:conf/aaai/WangH0SOC20] use item relations to enhance representation. yu-2020-predicting employ graph neural networks (GNNs) to model item-item relations between baskets and a self-attention mechanism to discern temporal dynamics. Some methods [DBLP:conf/sigir/BaiNZZDW18, DBLP:journals/corr/abs-2109-11654, DBLP:conf/dasfaa/LengYXX20, DBLP:conf/sigir/SunBDL0L20, DBLP:journals/jcst/WangZNG19] use auxiliary information, including product categories, amounts, prices, and explicit timestamps.
@@ 6.5.2 Conversational recommendation
The task of a conversational recommender system (CRS) is to provide recommendations to users through conversational interactions. CRSs are increasingly attracting attention [see, e.g., zhao2013interactive, ConversationRS-KDD16, yu2019visual, zou2020neural, mangili2020bayesian, sun2018conversational, lei20estimation, lei2020interactive, zhou2020improving, liu2020towards, zhou2020topicguided, zhang2020conversational, li2020seamlessly]. According to gao-2021-advances, the task of CRS is formally defined as follows:
@@
Building on advances in interactive recommendation [christakopoulou2018q, wang2017factorization, liu2020diversified], early studies on CRSs formulate the task as a specific application of task-oriented multi-turn dialogue systems (TDS) [DBLP:conf/ccks/LeLWWLJ18, dhingra-EtAl:2017:Long1, wen2017network, zhang2019memory]. Studies into CRSs follow one of two main types of strategy: attribute-aware and topic-guided.
@@
Attribute-aware CRSs aim to answer three main research questions: “whether to ask or recommend,” “which attributes to ask” or “which items to recommend.” Early work on attribute-aware CRSs obtains user preferences based on asking about items directly [zhao2013interactive, wang2018online, ConversationRS-KDD16, zou2020neural, DBLP:conf/aaai/VendrovLHB20], or asking attributes through a heuristic method [christakopoulou2018q, zhang2018towards, DBLP:conf/sigir/LuoYWS20]. There are two main kinds of attribute-aware CRSs solutions. One kind asks a fixed number of questions and makes a recommendation at the last turn [lei20estimation, lei2020interactive]; whereas the other predicts a specific turn to recommend items. Reinforcement learning strategies have successfully been applied to attribute-aware CRSs. liu2020towards and li2020seamlessly focus on cold-start users in conversational recommendation and extend bandit-based algorithms to balance the trade-off between exploration and exploitation. zou2022improving propose TSCR, a transformer-based sequential conversational recommendation method that captures the sequential dependencies in dialogues to enhance recommendation accuracy. deng2023unified propose a novel unified multi-goal conversational recommender system, named UniMIND, which unifies these goals into a single sequence-to-sequence (Seq2Seq) paradigm and employs prompt-based learning strategies to facilitate multi-task learning.
@@
Topic-guided CRSs interact with users through natural language conversations with fluent responses and precise recommendations [DBLP:conf/nips/LiKSMCP18, zhou2020topicguided, DBLP:conf/emnlp/ChenLZDCYT19, liu2020towards, zhou2020improving, DBLP:conf/emnlp/MaTH21, DBLP:conf/wsdm/ZhouZZWJ022]. Unlike attribute-aware CRSs, topic-guided CRSs focus on making recommendations using free text, which creates considerable flexibility to influence how a dialogue continues. External knowledge has been applied in topic-guided CRSs [DBLP:conf/emnlp/MaTH21, zhou2020improving, DBLP:conf/emnlp/ChenLZDCYT19]. DBLP:conf/emnlp/ChenLZDCYT19 integrate a recommendation system and a dialogue system via an end-to-end framework to bridge the gap between the two systems. DBLP:conf/nips/LiKSMCP18 use an auto-encoder for recommendation and a hierarchical RNN for response generation. zhou2020topicguided propose a topic-guided CRSs method that incorporates topic threads to enforce transitions actively toward a final recommendation. More recently, external knowledge graphs have been shown to be effective in improving the performance of topic-guided conversational recommendation systems. DBLP:conf/emnlp/ChenLZDCYT19 apply knowledge graphs to enhance the semantics of contextual items for recommendation. zhou2020improving incorporate both word-oriented and entity-oriented knowledge graphs. DBLP:conf/emnlp/MaTH21 perform tree-structured reasoning on a knowledge graph for recommendation. zhang2022analyzing focus on user reformulation behaviors to improve the robustness of conversational agents. ren2022variational explore user preferences in conversational recommendation and propose a variational reasoning mechanism to jointly track both short-term and long-term user behaviors. zhang2023variational present the first attempt to explicitly address the problem of dynamic reasoning over incomplete knowledge graphs. However, no study is capable of fusing recommendation and response generation in an end-to-end manner, which limits the potential for mutual reinforcement between these two tasks. Additionally, a lack of interpretability in current conversational recommendation system (CRS) models further hinders their ability to fully align with user needs. Models are typically trained on conversational recommendation datasets, but the assumption that the standard items and responses in these benchmark datasets are optimal leads to a tendency for CRSs to replicate the logic of the recommenders found in the data, rather than truly addressing the evolving needs of the users. This misalignment remains a significant challenge in advancing more user-centric and adaptable conversational recommendation systems.
@@
Although CRSs have many merits, their evaluation is still a thorny issue. Recent studies have evaluated CRSs either through offline evaluation or human evaluation [Lamel2000TheLA, li2015toward]. Offline evaluation evaluates a dialogue system based on test sets, whereas human evaluation reflects the overall performance of the agent through in-field experiments [Black2011SpokenDC, gilotte2018offline] or crowd-sourcing [zhou2020topicguided, DBLP:conf/nips/LiKSMCP18]. However, offline evaluation is often limited to single turn assessments, while human evaluation is intrusive, time-intensive, and is not scalable [zhao2019toward, siro-2022-understanding]. As an alternative, user simulators that mimic user behavior are able to provide broad insights to generate human-like conversations for assessing CRSs [afzali2023usersimcrs].